Consider the following short run production function L=variable input Q= output Q=10L-0.5L^2. Suppose that output can be sold for $10 per unit. Also assume that the firm can obtain as much of the variable input as it needs at $20 per unit.
A. Determine marginal revenue product function
B. determine marginal factor cost function
C. Determine the optimal value of L, given the objective is to maximize profits
1
Expert's answer
2014-12-08T10:56:01-0500
Q=10L-0.5L^2. P = $10 per unit, Pl = $20 per unit.
A. Marginal revenue product function is:
MRP = MR*MP = 10*Q' = 100 - 10L
B. Marginal factor cost function MC = Q' = 10 - L
C. The optimal value of L, given the objective is to maximize profits, is when MRP = 0, so 100 - 10L = 0, L = 10, so 10 workers should be hired to maximize profits.